A sample is counted and found to have 952 counts per minute. Seven minutes later it is measured again and has a count of 148 counts per minute. A background measurement gave 6 counts per minute. What is the half-life of the sample?
Answer #1

Activity of my time it a radioactive substance at is given by - N=N. At where a = decay Constant ont =N No - No at = lu(No) PPut No=982-6 = 946 Count / min N = 148-6=142 Count /min et = 7 min I 7x ln (2) = 2.56 min. = £ en Iu (94 946 1427 )